how to WRITE Y-1=2(X-3/2) 9N STANDARD FORM

Answer by NancyLam(35)

Y-1=2(X-3/2)
y-1=2x-6/2 mutiply
y-1=2x-3
y-2x=-3+1
y= 2x-2 ( slope intercept from)
STANDARD FORM:x+y=c, than you have
y-2x =-2
I hope it help